jordan is now three times as old as fred.In 5 years,jordan will be twice as old as fred will be then.What will thier ages be in 5 years?
Answer by Maths68(1474)

Let
Present age of Jordan = j
Present age of Fred = f
Jordan is now three times as old as Fred
j=3f..............(1)
In 5 years,
Jordan will be = j+5
Fred will be = f+5

Then
Jordan will be twice as old Fred
j+5=2(f+5)
j+5=2f+10
j=2f+10-5
j=2f+5..............(2)
Put the value of j from (1) to (2)
3f=2f+5
3f-2f=5
f=5

Put the value of f in (1)
j=3f
j=3(5)
j=15

Present age of Jordan = j = 15
Present age of Fred = f = 5


In five years their ages will be
Jordan = 15+5 = 20
Fred = 5+5 = 10